Week commencing 1 january 2018
Between Christmas and new year we visited the Grosmont workshop of the North York Moors Railway to examine our completed bogie axleboxes before accepting delivery to York. The boxes were measured and compared to our specification, and they will now be delivered to York.
While at Grosmont we visited our storage container and retrieved the boiler blanking plates, of which there are quite a number. Once back at York they were marked and listed so we know what we have. They will be sent to Llangollen for testing the boiler.
The machining of the piston valve heads is now well underway and we expect delivery of the ones machined from blank castings soon. Unfortunately, one of the old valve heads we had planned to machine to our new required dimensions was found to be dimensionally unsuitable when it was set up in the lathe, so this will have to be scrapped and a new casting ordered. As we also want to have a couple of spare castings in reserve, we decided that the last old valve head that we were going to re-machine should also be replaced with new, so four new castings have been ordered. The pattern has now been sent to the foundry.
The leading vacuum cylinder outside bearings were machined to finished dimensions over the Christmas break, along with the boring of the blow off cock. The lubricating plate for the air pump, which had been pre-drilled, received some milling to match the slotted holes on the face of the air pump. The outside of the plate will now be milled to shape.
The painting around the loco progresses. The coupled wheels have now received their first coat of top coat gloss, and they look really good. The axles of the bogie sets are now receiving their white gloss and further gloss coats have been given to the vacuum cylinders. The cladding sheets for under the middle cylinder received another coat of primer. Painting also continues around the loco, particularly at the dragbox end, when the pipe fitters are not in.
Pipe fitting progresses around the dragbox area. This week the damper shaft that runs across the frames in this area was fitted so that the pipe runs could be accurately placed. The injectors are now also in place, so their pipework can be routed. When pipes are being fitted we also work on the ashpan. This week wasted sections and some areas too warped to be straightened were cut out on the right side and front.
With the paint dried, the final assembly of the gravity sands operating gear was progressed with final pinning, while the cylinder drain cock operating gear received more cleaning off. New bushes are being made and fitted to the drain cock valves. These have to be accurately finished to reduce the steam leakage from the poppet valve stem to a minimum.
We have measured the distances between the rear fixed coupled wheel horns, and the measuring of the frames is now nearing completion . The alignment wire was taken down on Saturday and permission has been granted by our Chief Mechanical Engineer for the various rods and bars used for measuring to be removed from the frames. These dimensions will have to be processed and then final specifications can be issued for the machining of the coupled wheel axleboxes.
On the subject of the coupled wheel horns, the vertical horn stay bolting was examined in detail. The horn stays have previously been removed and closely inspected, which led to the weld repair of two of the horn stays. We don’t have any concerns regarding the stays’ fit, as they have been previously examined and are good. However, many of the nuts are showing signs of years of hammer examinations and flogging up. The measurements and observations taken will be considered and a repair procedure agreed with the CME.
Week commencing 8 January
We have been informed that the GSMR system requires a software upgrade, and that the body carrying this out requires the system to be powered up, preferably by the system batteries. As we have now removed the system from the tender this presented a bit of a problem: the batteries are charged with a special charger via the driver’s interface unit which is still on the front of the cab. However, we managed to release the connecting cables from the tender conduit and the batteries are now charged. We now await a date in the near future for the software upgrade.
The loco brake beams that are being contractor repaired were inspected this week. The worn surfaces have been rebuilt and they are now being put up on a lathe to check for straightness. Where required they are being heated and straightened and the end shafts are being machined.
The finished bored outside leading vacuum cylinder bearings have been tried in position and the alignment is correct. The centre bracket is fitted with new bolting and now fitted grease nipples have been fitted to the bearings in an effort to maximise the life of the cylinder trunnions.
The cylinder drain cock overhaul continues, with the poppet valve spindle bushes now all machined and fitted. The integral relief valves also required skimming and have now all been lapped in. Meanwhile, the inspection of the operating linkage has been completed and it has been refitted it to the loco. New pins will be fitted as necessary from blanks machined by volunteers from the engineering team.
The mammoth job of painting the wheels continues with the final coat of black gloss. Painting of the bogie wheels continues with gloss being applied to axles and wheels.
Piping around the cab end frames continues. With pipe runs in this area dependent upon where they go into the cab, it was decided that it was time to put the cab floor plate in, as this dictates where pipes come up into the cab. The plate was retrieved and moved to a space near the cab and is now in the process of being prepared for painting. The frame top where the plate will be fitted is also being painted ready to receive the plate.
The last of the hard pipes to the trailing vacuum cylinder have now been fitted and the flexi pipes to all the cylinders have been retrieved from store and inspected. The trailing vacuum cylinder was upturned this week and the bottom face glossed.
More wasted plate was cut from the ashpan and new plate trimmed to size and prepped for welding in place. We plan to have the welder in next week.
Now that the frame horns have been measured, attention is turning to the coupling rods. These were positioned next to the frames this week and cleaned down. The remaining oil was removed from the reservoirs and has been disposed of. The coupling rod bushes were then measured and these measurements have been sent to our CME for consideration.
The upper slide bars are now being examined for wear. The gap between top and bottom bars were measured before they were removed, but this does not give us a measurement for flatness. So the bars are being put up on the surface table in the workshop and the flatness measured. Two of the upper slide bars were measured and the third is still to be done. The table surface was prepared and its flatness was checked with a straight edge.
One of the operating cams in the vacuum ejector was found to be cracked. It is very small, but we have decided to make a new cam. When the can was measured we found that the centre hole where it fits onto the operating shaft is distorted, so this week the shaft was removed and the new cam will be made to fit the shaft.
The superheater face on which the steam pipes are connected was cleaned off. It initially didn’t look too bad, though we did expect to spend some time preparing it for sealing—but when the smooth, hard outer layer was wire brushed, it exposed some quite deep corrosion and perhaps evidence of previous weld repairs. Concerned that this may present a leak path in future, we decided to more thoroughly repair this face. As a start the face will have to be cleared of studs. Again this is unfortunate because the condition of the studs is good and we had hoped to reuse then. The studs were heated and they were mostly easily removed by double nutting, with a few obstinate ones removed by the Stillsons. Most of the stud holes have been cleaned out with a tap.
This is the 24th update. You can read the previous instalments here.